A motor pumps 1000 litres of water in one second upto a height of 10m. If the efficiency of motor is 70%, find the power required to run it ? 

Grade:10

1 Answers

Adarsh
9 years ago
Let the power required to run it be P
Then 70% of P = mgh = ρVgh
where m – mass of water, g – acceleration due to gravity, h – height, ρ – density of water, V – volume of water
So P = ρVgh/70% = 1000*1000*9.8*10 / (70/100)=1.4*108
 
Please approve if my answer is correct.